FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE E^{2}Π–X^{2}Σ^{+} TRANSITION OF CaH AND CaD

\underline{R. S. RAM}, K. TERESZCHUK and P. F. BERNATH, Department of Chemistry, University of York, Heslington, York YO10 5DD, UK; I. E. GORDON, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138, USA; K. A.WALKER, Department of Physics, University of Toronto, Toronto, Ont., M5S 1A7, Canada.

The emission spectra of CaH and CaD have been recorded at high resolution using a Fourier transform spectrometer and bands belonging to the E^{2}Π–X^{2}Σ^{+} transition have been measured in the 20100–20700 cm^{−1} region. A rotational analysis of 0–0 and 1–1 bands of both the isotopologues has been carried out. The present measurements have been combined with the previously available pure rotation and vibration-rotation data to provide improved spectroscopic constants for the E^{2}Π state. The constants ΔG_{1/2} = 1199.8810(32) cm^{−1}, B_{e}=4.344659(45) cm^{−1}, α_{e}=0.121869(88) cm^{−1}, r_{e}=1.986718 Å for CaH, and ΔG_{1/2}=868.7438(46) cm^{−1}, B_{e}=2.212496(51) cm^{−1}, α_{e}=0.036509(97) cm^{−1}, r_{e}=1.993396(23) Å for CaD have been determined. An analysis of the corresponding transitions of SrH and SrD in the 18600–19300 cm^{−1} region will also be reported.